OpenStudy (anonymous):

Determine the approximate value of g, at height D above the earth's surface, where D is the diameter of the earth.

OpenStudy (anonymous):

Ok.. can you start somewhere?

OpenStudy (anonymous):

Vorbit \[=\sqrt{rg}\]

OpenStudy (anonymous):

g\[= (V _{orbit})^{2}/r\]

OpenStudy (anonymous):

\[\frac{ 7900^{2} }{ 6.37 \times 10^{6}+h }\]

OpenStudy (anonymous):

question is, what is h? is it 3 x r or 2x r. i would like to think that it is 2xr since the question says earth's surface, not earth's atmosphere, or is it the same thing?

OpenStudy (unknownrandom):

I'm kind of thinking 2xr also. What level physics is this?

OpenStudy (anonymous):

first year university, A LEVEL and or second year in college

OpenStudy (unknownrandom):

Okay, my input might not be valuable considering I'm in a high school physics class.

OpenStudy (anonymous):

all is valuable. it's high school final year stuff. ok , if it is D from the surface of the earth, then , THE Denominator or radius should be 2r.

OpenStudy (anonymous):

@Potatoes.ramu

OpenStudy (anonymous):

h is 2 times the radius = Diameter.. which is what is mentioned in your question!

OpenStudy (anonymous):

i thought as much, but for some reason, i got confused. but thank you @Mashy

OpenStudy (prasadrajupvmv):

g'=g(1-(2h/d))

OpenStudy (anonymous):

Wooow @prasadrajupvmv . am dead lost!

OpenStudy (prasadrajupvmv):

what happend ? @MERTICH

OpenStudy (anonymous):

g'?

OpenStudy (prasadrajupvmv):

the new g value at height h